OpenStudy (anonymous):

A ball is dropped from a height of 20 m above the surface of water in a lake. The refractive index of water is 4/3. A fish inside the lake, in the line of fall of the ball, is looking at the ball. At an instant, When the ball is 12.8 m above the water surface, the fish sees the speed of ball as

OpenStudy (anonymous):

@Mashy dude.. :D

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Ek baat bataun?

Parth (parthkohli):

Mujhe Physics nahi aati... tag mat karo >_>

OpenStudy (anonymous):

Tereko sab aata hai. Mereko pta hai. :)

Parth (parthkohli):

Nahi aata na!

OpenStudy (anonymous):

Ok. :/

OpenStudy (anonymous):

hmmm lemme think!

OpenStudy (anonymous):

Ho gya. :D

OpenStudy (anonymous):

yea .. i got it too! :D

OpenStudy (anonymous):

How did you do it btw. Just to see if your method is easier.

OpenStudy (anonymous):

no no first check whether my method is correct :P.. lemme type

OpenStudy (anonymous):

First find the height of the ball as seen by the fish when it is 20m above and when it is 12.8m above .. using normal shift formula then find the difference .. that gives the apparent displacement of the ball as seen by the fish then use vsquare = 2gh where h is the calculated apparent displacement?

OpenStudy (anonymous):

Wait. I did that too. I think i solved it wrong.

OpenStudy (anonymous):

m high on pizza.. eating my 7th piece.. so cut me some slack if i do some mistake :D

OpenStudy (anonymous):

x-ap=80/3?

OpenStudy (anonymous):

I mean x(ap). Not x (minus) ap . :P

OpenStudy (anonymous):

And no i didnt get it. v=sqrt2gh=12m/s Then using the shift formula. x(image of the ball)=4/3 x(ball) .. What next?

OpenStudy (anonymous):

Hmm. So differntiating both sides wrt time gives me the answer. I am not sure if that is the correct approach.

OpenStudy (anonymous):

wait.. aph = 20* 3/4 = 15m and aph2 = 12.8*3/4= 9.6 hence displacement = 15 - 9.6 = 5.4 hence v square = 2*9.8*5.4 that doesn't give the answer?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

ahh.. so m sure it sees that the ball has travelled that much distance .. hmmmm

OpenStudy (anonymous):

What? :o

OpenStudy (anonymous):

arre.. the fish sees the ball to be having that much displacement only na??? 5.4m displacement?! .. but i think it the acceleration no longer remains g.. yea!!.. things get screwed up..

OpenStudy (anonymous):

Why doesnt acc remain g?

OpenStudy (anonymous):

cause everything is normal shifted.. and the normal shift depends on the depth of the object as well.. so as the object keeps coming closer and closer, the Normal shift keeps decreasing... hence no longer is it accelerating at g .. i dun even know if the acceleration remains constant :P

OpenStudy (anonymous):

Ok so differentiating h(ap)=h(actual)*myu, seems logical?

OpenStudy (anonymous):

wrt time...

OpenStudy (anonymous):

hmm.. yea..so all you do is find the actual velocity and multiply by myu? and besides i did it wrong earlier anyways :D

OpenStudy (anonymous):

Haha.. Another question coming up. i'll upload. Wait.

OpenStudy (anonymous):

wait.. so is taht it??

OpenStudy (anonymous):

Yup. That gives me the answer.

OpenStudy (anonymous):

its definitely not 1 and 4 for sure

OpenStudy (anonymous):

it can't be 3.. also!!

OpenStudy (anonymous):

i t has to be 2

OpenStudy (anonymous):

Yes, the critical angle thing. Why not 3?

OpenStudy (anonymous):

because in three.. the transmission is at 100%.. if that was the case.. the reflected ray should have been at zero percent but its not.. so its impossible.. !!

OpenStudy (anonymous):

initially i mean!!

OpenStudy (anonymous):

:O I get it... Thank you :D

OpenStudy (anonymous):

ok .. so the medal is for THIS question.. and not the previous one.. :D

OpenStudy (anonymous):

Haha yeah. :D

OpenStudy (anonymous):

More questions tomorrow. :P

OpenStudy (anonymous):

are you into dbz??

OpenStudy (anonymous):

that is for the earlier qusetion :D

OpenStudy (anonymous):

dbz? And thank you.

OpenStudy (anonymous):

dragon ball z?

OpenStudy (anonymous):

Oh, i used to love it. But not now. I dont like watching television anyway.